The sum of the
measure of the interior angles in a
triangle is 180°.The measure of the
second angle is 15° more than the
measure of the first angle.The measure
of the third angle is 45° more
than the measure of the first angle.
Find the measures of the interior angles in the
triangle.

Answer by JBarnum(2146)

The sum of the
measure of the interior angles in a
triangle is 180°.The measure of the
second angle is 15° more than the
measure of the first angle.The measure
of the third angle is 45° more
than the measure of the first angle.
Find the measures of the interior angles in the
triangle.
ok take line by line and make up variables im going to do < A,B,C
"The sum of the measure of the interior angles in a triangle is 180°"
A+B+C=180
1st angle=A, 2nd angle=B, 3rd angle=C
"The measure of the second angle is 15° more than the measure of the first angle" B=A+15
"The measure of the third angle is 45° more than the measure of the first angle" C=A+45
"Find the measures of the interior angles in the triangle." Find A,B,C
rewrite your equations
A+B+C=180 B=A+15 C=A+45
since we know what B and C equal, I would suggest solving for A first.
A+B+C=180
A+(A+15)+(A+45)=180
3A+60=180
3A=120
A=40 now plug into the other 2 equations
B=(40)+15=55
C=(40)+45=85
So A=40 B=55 and C=85